# Taming the Untamed: Graph-Based Knowledge Retrieval and Reasoning for MLLMs to Conquer the Unknown

### v2

@@ Abstract
The real value of knowledge lies not just in its accumulation, but in its potential to be harnessed effectively to conquer the unknown. Although recent multimodal large language models (MLLMs) exhibit impressing multimodal capabilities, they often fail in rarely encountered domain-specific tasks due to limited relevant knowledge. To explore this, we adopt visual game cognition as a testbed and select “Monster Hunter: World” as the target to construct a multimodal knowledge graph (MH-MMKG), which incorporates multi-modalities and intricate entity relations. We also design a series of challenging queries based on MH-MMKG to evaluate the models’ ability for complex knowledge retrieval and reasoning. Furthermore, we propose a multi-agent retriever that enables a model to autonomously search relevant knowledge without additional training. Experimental results show that our approach significantly enhances the performance of MLLMs, providing a new perspective on multimodal knowledge-augmented reasoning and laying a solid foundation for future research. [^1] The dataset and code will be released soon.
@@ 2.2 Multimodal Knowledge Graph
Notwithstanding these considerable achievements, current works primarily focus on integrating text and image modalities [38] via web data, which is already embedded in the knowledge of MLLMs. The exploration of additional modalities, such as audio and video, as well as practical applications of recent MLLMs for complex real-world tasks, remains limited. Our meticulously crafted MH-MMKG incorporates multiple modalities alongside complex relations as a knowledge base for a domain-specific task.
@@ 3.1 MH-MMKG
Let $\mathcal{G}=(\mathcal{E},\mathcal{V},\mathcal{R})$ denote our whole KG, where $\mathcal{E}$ is the set of entities, $\mathcal{V}$ is the set of edges, and $\mathcal{R}$ is the set of relations. $\mathcal{E}$ consists of two types of entities: 1) monsters in $\mathcal{E}_{\text{o}}$ and 2) other entities in $\mathcal{E}_{\text{a}}$. An edge $v=(e,r,e^{\prime})\in\mathcal{V}$ represents $e\in\mathcal{E}$ has relation $r\in\mathcal{R}$ with $e^{\prime}\in\mathcal{E}$. The attribute associated with an entity $e$ is denoted as $A(e)=(c,u)$, which comprises a video clip $c$ and/or textual context $u$. A video clip describes the entity $e$, recorded in 60 fps and 4k. As current MLLMs are not able to handle videos directly, the dataset also offers human-selected keyframes (no more than 10 frames) as well as human-written captions about $c$ [^3] The knowledgeable players who curated the dataset also selected the keyframes and wrote the captions. The textual context $u$ provides knowledge about the entity but is not included in $c$. We denote a set of all attribute as $\mathcal{A}=\{A(e)|e\in\mathcal{E}\}$.
@@
I: Individual Information Retrieve the textual information of a monster, e.g., nick name, habitat, and skill mechanics. 21
II: Attack Recognition Recognize the about to, ongoing or finished attack action of a monster. 97
III: Combo Premonition Predict upcoming attack sequences based on the monster’s current action or previous action track. 36
IV: Condition Awareness Detect status on monsters or surrounding environment, e.g., whether angry, terrain or phase changes, to anticipate future battle. 30
V: Proc Effect Insight Analyze the effects such as environment and attack on monster status and movement patterns. 40
VI: Cross Monster Analysis Compare attack patterns or behaviors across different monsters to optimize hunting strategies. 14
@@ 3.2 MH Benchmark
[…]
@@ 4.1 Multi-agents Retriever
$\mathcal{O}_{t+1}=\mathcal{E}^{\prime}_{t+1}\backslash\mathcal{E}^{\prime}_{t},$ (9)
[…]
@@ 5 Results
GPT-4o [1] .3361 .4033 .9076 .7353 .4545 .6667 .4412 .2500 .4406 .5420 .2715 .5881
GPT-4o mini [1] .3193 .3697 .8361 .6260 .2821 .4952 .4243 .1643 .4434 .3781 .1522 .2835
Claude 3.7 Sonnet [2] .2941 .3403 .9245 .7017 .5681 .6254 .3950 .3002 .3591 .4538 .3069 .4624
Claude 3.5 Sonnet [2] .3151 .3571 .8823 .7142 .5653 .6318 .4160 .3335 .4252 .4075 .2419 .3591
Claude 3.5 Haiku [2] .2605 .3151 .8823 .6344 .3334 .4463 .3781 .1945 .2720 .3907 .1669 .3161
Gemini 2.0 Flash [47] .1890 .2689 .8361 .6596 .3531 .6187 .3823 .1838 .4070 .4159 .1969 .4367
Gemini 1.5 Pro [47] .2394 .2521 .8697 .6932 .4939 .6149 .2899 .1634 .2183 .4075 .2156 .4406
@@
InternVL2.5-78B-MPO [12] .1596 .3487 .8949 .6596 .4233 .5076 .3781 .2027 .3007 .3361 .1591 .3235
Qwen2.5-VL-72B [3] .1512 .2731 .8613 .5084 .3284 .3783 .3025 .1622 .1724 .2394 .1463 .1839
Ovis2-16B [36] .1974 .3613 .8655 .5966 .3286 .4406 .2521 .1141 .2817 .2863 .1086 .3122
MiniCPM-o-2.6 [53] .2058 .2857 .8025 .4495 .1761 .2534 .2058 .0590 .1609 .2643 .0594 .2268
DeepSeek-VL2-Small [52] .1344 .2226 .7100 .3058 .0762 .0951 .1890 .0603 .0484 .1512 .0363 .0373
@@ 5.1 Experimental Settings
[…]
@@ 5.2 Performance of MLLMs
As shown in Table 3, with the vanilla setting, all methods hardly predicted correct answers. GPT-4o achieves the best performance with an accuracy of 0.34, whereas the knowledgeable human attains 0.53. The closed-source models generally outperform open-source ones, implying that the closed-source models have richer built-in knowledge about the game. With the vanilla+ setting, which allows human captions $s$ instead of images $d$, most models’ performance improved, which means that MH Benchmark requires comprehension of visual context. In the knowledgeable experiment, the models achieved an accuracy of roughly 0.9. We thus argue that MH-MMKG provides sufficient knowledge to answer the questions in MH Benchmark.
@@
The perceptive, unaided-offline, and unaided-online settings evaluate a model’s ability to retrieve relevant knowledge In the perceptive setting, which uses human captions for both queries and knowledge, all models demonstrated a strong capacity to identify relevant knowledge, significantly improving performance compared to vanilla+. Additionally, the models exhibit high knowledge consistency. In the unaided-offline setting, all models performed worse than in the perceptive setting, suggesting that visual perception plays a crucial role in knowledge retrieval. Nevertheless, they are still still better than the vanilla setting, demonstrating the benefits of knowledge retrieval. The unaided-online setting further challenges the models’ visual perception to generate captions with richer context (i.e., $Q$ and the knowledge on the path in $\hat{\mathcal{I}}$) compared to the unaided-offline. Our results show that only some models surpassed the offline variant (highlighted in green), yet the improvement is evident—–GPT-4o improves by 0.1, and Claude 3.7 Sonnet by 0.06. This suggests that knowing what to do enhances a model’s visual perception, strengthening its planning ability, and ultimately results in improved reasoning ability. Additionally, we find that recall is consistently higher than precision, suggesting that models tend to retrieve more knowledge. On the contrary, humans showed higher precision (though both precision and recall are high compared to MLLMs).
@@
We also analyzed the performance differences in the unaided-online setting across six sub-tasks over GPT-4o, Claude 3.7 Sonnet, and Gemini 1.5 Pro, as illustrated in Figure 4. GPT-4o achieves the best performance in all sub-tasks except II. It is particularly strong in sub-tasks IV and V but falls short in sub-tasks II and III, which are more reliant on fine-grained visual perception. Additionally, sub-task I is the simplest, involving only a single relationship $v$, for which the models, except for Gemini, generally perform well. These results suggest that visual perception remains a key challenge for MH benchmark.
@@
Figure 5 presents six examples (one per sub-task) in the vanilla and unaided-online settings, where all predictions are by GPT-4o. In the retrieved path panes, the green paths are both in $\mathcal{I}$ and $\hat{\mathcal{I}}$ (i.e., correctly retrieved paths), the dotted green paths are in $\mathcal{I}$ but not in $\hat{\mathcal{I}}$, and the black paths are in $\hat{\mathcal{I}}$ but not in $\mathcal{I}$. The key relationship or essential information for deriving the correct answer is highlighted in orange. We report the precision and recall metrics for each example. GPT-4o retrieves many knowledge paths, capturing all relevant knowledge for prediction, though precision is low. In Figure 5, we can see that GPT-4o actually retrieved many paths for sub-tasks III and VI. Interestingly, for challenging cases like sub-tasks II and IV, GPT-4o successfully recognized detailed visual concepts (e.g., the monster flying backward and the monster sticking to the wall) and generated plausible captions, which require fine-grained visual perception.
@@ 5.3 Analysis of Factors Affecting Performance
Impact of Captioning Performance. Table 4 compares performance scores of captioning, where the metric similarity (Sim.) measures the similarity between generated and human captions using GPT-4o (see supplementary materials for details). The table also summarizes the corresponding reasoning accuracy and knowledge consistency scores. The results indicate that the unaided-online setting consistently gives higher similarity scores than the unaided-offline, suggesting that awareness of query $Q$ and retrieval process consistently enhances captioning performance. Also, the similarity scores positively correlated with the reasoning accuracy scores. We also evaluated the unaided-offline variant on GPT-4o but its offline captions are replaced with ones by video models InternVideo2.5 [50] and VideoChat-Flash [32]. The results showed lower performance than the original GPT-4o across all metrics, indicating that current video models have limited capability in visual comprehension for MH Benchmark.
@@
GPT-4o [1] .4412 .2500 .4406 .2715
GPT-4o [1] .5462 .2715 5881 .2913
Claude 3.7 Sonnet [2] .3950 .3002 .3591 .2649
Claude 3.7 Sonnet [2] .4538 .3069 .4624 .3162
Gemini 1.5 Pro [47] .2899 .1634 .2183 .1643
Gemini 1.5 Pro [47] .4075 .2156 .4406 .1712
InternVideo2.5 [50] .3697 .1960 .2959 .1669
VideoChat-Flash [32] .3445 .2135 .2863 .1421
Keyframe Selection. Due to the limited number of input tokens, current MLLMs cannot handle videos, and keyframe sampling is necessary. MH-MMKG provides human-selected keyframes to facilitate the evaluation on this point. To show the difference between human-selected keyframes and typically-adopted equal-interval sampling, we compare them with GPT-4o and Claude 3.7. Table 5 summarizes the results, where sampling is done in two frames per second, and the maximum number of frames is capped at ten frames. The keyframe selection strategy does impact the performance: Human-selected keyframes seem to provide more informative visual cues, though the difference is not substantial.
@@
Settings Model Acc. Pre. Rec. Acc. Pre. Rec.
Human GPT-4o .4412 .2500 .4406 .5462 .2715 .5881
Sampling GPT-4o .4117 .2363 .4061 .5037 .2494 .5368
Human Claude 3.7 .3950 .3002 .3591 .4538 .3069 .4624
Sampling Claude 3.7 .3571 .2798 .3248 .4033 .2739 .3749
@@
The Number of Paths Used in Reasoning. For the same reason, the number of the paths in $\hat{\mathcal{I}}$ used for reasoning is limited. All evaluations so far used 5 random paths, [^6] Our baseline uses the first 5 paths found in the retrieval. A shorter path is thus preferred. though this number can change the performance. Figure 6(a) shows the relationship between the number of paths and the reasoning accuracy. GPT-4o gives the optimal performance when 5 paths are used. We also show in the figure the number of queries $Q$ for which at least a certain number of paths are retrieved. GPT-4o tends to find more paths as its knowledge (the recall is higher), while Claude is more conservative in retrieval (the precision is higher).

### v3

@@ Abstract
The real value of knowledge lies not just in its accumulation, but in its potential to be harnessed effectively to conquer the unknown. Although recent multimodal large language models (MLLMs) exhibit impressing multimodal capabilities, they often fail in rarely encountered domain-specific tasks due to limited relevant knowledge. To explore this, we adopt visual game cognition as a testbed and select “Monster Hunter: World” as the target to construct a multimodal knowledge graph (MH-MMKG), which incorporates multi-modalities and intricate entity relations. We also design a series of challenging queries based on MH-MMKG to evaluate the models’ ability for complex knowledge retrieval and reasoning. Furthermore, we propose a multi-agent retriever that enables a model to autonomously search relevant knowledge without additional training. Experimental results show that our approach significantly enhances the performance of MLLMs, providing a new perspective on multimodal knowledge-augmented reasoning and laying a solid foundation for future research. [^1] The dataset and code at https://github.com/wbw520/MH-MMKG.
@@ 2.2 Multimodal Knowledge Graph
Notwithstanding these considerable achievements, current works primarily focus on integrating text and image modalities [38] via web data, which may already embedded in the knowledge of MLLMs. The exploration of additional modalities, such as audio and video, as well as practical applications of recent MLLMs for complex real-world tasks, remains limited. Our meticulously crafted MH-MMKG incorporates multiple modalities alongside complex relations as a knowledge base for a domain-specific task.
@@ 3.1 MH-MMKG
Let $\mathcal{G}=(\mathcal{E},\mathcal{V},\mathcal{R})$ denote our whole KG, where $\mathcal{E}$ is the set of entities, $\mathcal{V}$ is the set of edges, and $\mathcal{R}$ is the set of relations. $\mathcal{E}$ consists of two types of entities: 1) monsters in $\mathcal{E}_{\text{o}}$ and 2) other entities in $\mathcal{E}_{\text{a}}$. An edge $v=(e,r,e^{\prime})\in\mathcal{V}$ represents $e\in\mathcal{E}$ has relation $r\in\mathcal{R}$ with $e^{\prime}\in\mathcal{E}$. The attribute associated with an entity $e$ is denoted as $A(e)=(c,u)$, which comprises a video clip $c$ and/or textual context $u$. A video clip describes the entity $e$, recorded in 60 fps and 4k. As current MLLMs are not able to handle videos directly, the dataset also offers human-selected keyframes (no more than 10 frames) as well as human-written captions about cc. [^3] The knowledgeable players who curated the dataset also selected the keyframes and wrote the captions. The textual context $u$ provides knowledge about the entity but is not included in $c$. We denote a set of all attribute as $\mathcal{A}=\{A(e)|e\in\mathcal{E}\}$.
@@
I: Individual Information Retrieve the textual information of a monster, e.g., nick name, habitat, and skill mechanics. 24
II: Attack Recognition Recognize the about to, ongoing or finished attack action of a monster. 109
III: Combo Premonition Predict upcoming attack sequences based on the monster’s current action or previous action track. 28
IV: Condition Awareness Detect status on monsters or surrounding environment, e.g., whether angry, terrain or phase changes, to anticipate future battle. 29
V: Proc Effect Insight Analyze the effects such as environment and attack on monster status and movement patterns. 35
VI: Cross Monster Analysis Compare attack patterns or behaviors across different monsters to optimize hunting strategies. 13
@@ 3.2 MH Benchmark
$\textit{Rathian}\xrightarrow{\text{attack of}}\textit{Triple Rush}\xrightarrow{\text{continues}}\textit{Bite},$ (1)
@@ 4.1 Multi-agents Retriever
where $U$ again is an MLLM with a carefully designed prompt. $o(e)=\textit{Yes}$ if the knowledge is good enough and no more expansion is needed; otherwise, $o(e)=\textit{No}$ and continues the expansion  [^5] UU also output a caption for $c$ of the entity $e$ in unaided-online setting..
@@
where the operator $\backslash$” represents set subtraction.
@@ 5 Results
GPT-4o [1] .3122 .3924 .8565 .7383 .5061 .7046 .4050 .2595 .4416 .5105 .2756 .5625
GPT-4o mini [1] .3218 .4135 .8481 .6877 .2963 .5450 .4514 .2059 .5028 .3544 .1626 .3009
Claude 3.7 Sonnet [2] .2827 .3375 .8987 .7004 .5817 .6322 .3628 .2775 .3270 .4388 .2911 .4029
Claude 3.5 Sonnet [2] .2869 .3755 .8776 .7215 .5922 .6800 .3966 .3215 .4008 .3966 .2330 .3270
Claude 3.5 Haiku [2] .2356 .3206 .8823 .6455 .3739 .5007 .3544 .2002 .3164 .3670 .1735 .3361
Gemini 2.0 Flash [47] .1983 .2995 .8438 .6919 .3507 .6146 .3839 .1703 .4092 .3713 .1515 .3663
Gemini 1.5 Pro [47] .2194 .2700 .8438 .6962 .4761 .6033 .3164 .1615 .2194 .4050 .2122 .4585
@@
InternVL2.5-78B-MPO [12] .1603 .2616 .8649 .5991 .4198 .5428 .2700 .1729 .2250 .3080 .1556 .2378
Qwen2.5-VL-72B [3] .1476 .2616 .8734 .6244 .4602 .4908 .3206 .2139 .2383 .3164 .1615 .1814
Ovis2-16B [36] .1645 .2573 .8902 .7046 .5407 .5949 .2869 .1963 .2383 .3459 .1853 .2878
MiniCPM-o-2.6 [53] .1139 .2405 .8312 .4683 .3183 .4001 .2194 .1311 .2376 .1687 .0189 .0210
DeepSeek-VL2-Small [52] .1139 .2362 .6455 .3586 .1419 .1708 .1814 .0400 .0759 .1181 .0042 .0042
@@ 5.1 Experimental Settings
Evaluation Metrics. We use accuracy (Acc.) to measure the correctness of predicted answers $\hat{y}$ compared to the ground-truth answer $y$. Since the questions in our benchmark follow an open-ended question-answering format, we employ GPT-4o as a judge with few-shot samples [59]. Knowledge consistency is also used to calculate the consistency between $\mathcal$$\hat${I}} and $\mathcal{I}$ using precision (Pre.) and recall (Rec.), demonstrating the efficiency in searching relevant knowledge. Human evaluation for GPT-4o as a judge and metric definition are shown in the supplementary material.
@@ 5.2 Performance of MLLMs
As shown in Table 3, with the vanilla setting, all methods hardly predicted correct answers. GPT-4o achieves the best performance with an accuracy of 0.31, whereas the knowledgeable human attains 0.53. The closed-source models generally outperform open-source ones, implying that the closed-source models have richer built-in knowledge about the game. With the vanilla+ setting, which allows human captions $s$ instead of images $d$, most models’ performance improved, which means that MH Benchmark requires comprehension of visual context. In the knowledgeable experiment, the models achieved an accuracy of roughly 0.9. We thus argue that MH-MMKG provides sufficient knowledge to answer the questions in MH Benchmark.
@@
The perceptive, unaided-offline, and unaided-online settings evaluate a model’s ability to retrieve relevant knowledge In the perceptive setting, which uses human captions for both queries and knowledge, all models demonstrated a strong capacity to identify relevant knowledge, significantly improving performance compared to vanilla+. Additionally, the models exhibit high knowledge consistency. In the unaided-offline setting, all models performed worse than in the perceptive setting, suggesting that visual perception plays a crucial role in knowledge retrieval. Nevertheless, they are still still better than the vanilla setting, demonstrating the benefits of knowledge retrieval. The unaided-online setting further challenges the models’ visual perception to generate captions with richer context (i.e., $Q$ and the knowledge on the path in $\hat{\mathcal{I}}$) compared to the unaided-offline. Our results show that only some models surpassed the offline variant (highlighted in green), yet the improvement is evident–GPT-4o improves by 0.1, and Claude 3.7 Sonnet by 0.07. This suggests that knowing what to do enhances a model’s visual perception, strengthening its planning ability, and ultimately results in improved reasoning ability. Additionally, we find that recall is consistently higher than precision, suggesting that models tend to retrieve more knowledge. On the contrary, humans showed higher precision (though both precision and recall are high compared to MLLMs).
@@
We also analyzed the performance differences in the unaided-online setting across six sub-tasks over GPT-4o, Claude 3.7 Sonnet, and Gemini 1.5 Pro, as illustrated in Figure 4. GPT-4o achieves the best performance in all sub-tasks except I and VI. It is particularly strong in sub-tasks IV and V, which are more reliant on fine-grained visual perception. Additionally, sub-task I is the simplest, involving only a single relationship $v$, for which the models generally perform well. For sub-task VI, although the accuracy is high, models fail to find the correct path (model uses its inherent knowledge for response). These results suggest that visual perception remains a key challenge for MH benchmark.
@@
Figure 5 presents six examples (one per sub-task) in the vanilla and unaided-online settings, where all predictions are by GPT-4o. In the retrieved path panes, the green paths are both in $\mathcal{I}$ and $\hat{\mathcal{I}}$ (i.e., correctly retrieved paths), the dotted green paths are in $\mathcal{I}$ but not in $\hat{\mathcal{I}}$, and the black paths are in $\hat{\mathcal{I}}$ but not in $\mathcal{I}$. The key relationship or essential information for deriving the correct answer is highlighted in orange. We report the precision and recall metrics for each example. GPT-4o retrieves many knowledge paths, capturing all relevant knowledge for prediction, though precision is low. Interestingly, for challenging cases like sub-tasks II, GPT-4o successfully recognized detailed visual concepts (e.g., the monster flying backward and the monster sticking to the wall) and generated plausible captions, which require fine-grained visual perception.
@@ 5.3 Analysis of Factors Affecting Performance
Impact of Captioning Ability. Table 4 compares performance of captioning, where the metric similarity (Sim.) measures the similarity between generated and human captions using GPT-4o (see supplementary for details). The table also summarizes the accuracy and knowledge consistency scores. The results indicate that the unaided-online setting consistently gives higher similarity scores than the unaided-offline, suggesting that awareness of query $Q$ and retrieval process consistently enhances captioning performance. Also, the similarity scores positively correlated with the reasoning accuracy scores. We also evaluated the unaided-offline variant on GPT-4o but its offline captions are replaced with ones by video models InternVideo2.5 [50] and VideoChat-Flash [32]. The results showed lower performance than the original GPT-4o across all metrics, indicating that current video models have limited capability in visual comprehension for MH Benchmark.
@@
GPT-4o [1] .4050 .2595 .4416 .2806
GPT-4o [1] .5105 .2756 .5625 .2948
Claude 3.7 Sonnet [2] .3628 .2775 .3270 .2776
Claude 3.7 Sonnet [2] .4388 .2911 .4029 .3208
Gemini 1.5 Pro [47] .3164 .1615 .2194 .1608
Gemini 1.5 Pro [47] .4050 .2122 .4585 .1746
InternVideo2.5 [50] .3697 .1960 .2959 .0525
VideoChat-Flash [32] .3445 .2135 .2863 .0644
Keyframe Selection. Due to the limited number of input tokens, current MLLMs cannot handle videos, and keyframe sampling is necessary. MH-MMKG provides human-selected keyframes to facilitate the evaluation on this point. To show the difference between human-selected keyframes and typically-adopted equal-interval sampling, Table 5 summarizes the results on GPT-4o, where sampling is done in two frames per second, and the maximum number of frames is capped at ten frames. The keyframe selection strategy does impact the performance: Human-selected keyframes seem to provide more informative visual cues, though the difference is not substantial.
@@
Settings Acc. Pre. Rec. Acc. Pre. Rec.
Human .4050 .2595 .4416 .5105 .2756 .5625
Sampling .3725 .2199 .3893 .4840 .2388 .5254
@@
The Number of Paths Used in Reasoning. In our experiments, the number of the paths in $\hat{\mathcal{I}}$ used for reasoning is limited. All evaluations so far used 5 paths, though this number can change the performance  [^6] Our baseline uses the first 5 paths. A shorter path is thus preferred.. Figure 6(a) shows the relationship between the number of paths and the reasoning accuracy. GPT-4o gives the optimal performance when 5 paths are used. We also show in the figure the number of queries $Q$ for which at least a certain number of paths are retrieved. GPT-4o tends to find more paths as its knowledge (the recall is higher), while Claude 3.7 is more conservative in retrieval (the precision is higher).